# Uncertainty, von Neumann entropy and squeezing in bipartite state of two-level atoms

### v1

@@ I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
We find that for the quantum state in Eq. (12),
$\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1+2AD+2BC\Big{)},$ (35)
$\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle=A^{2}+D^{2},$ (36)
$\langle\psi_{ab}|\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}|\psi_{ab}\rangle=(A-D)(B+C).$
Therefore, using Eqs. (25), (32), (34) to (I.2), we get
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}^{2}|\psi_{ab}\rangle=\frac{1}{2}\big{(}A+D\big{)}^{2}$ (38)
So, using Eqs. (24) and (38), we get
@@
Similarly,
@@
These are the coordinate frame independent inherent uncertainties of the system.
We now proceed to calculate similar uncertainties in the atomic operators $\hat{J}_{a_{x^{\prime}}}$, $\hat{J}_{a_{y^{\prime}}}$, $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ of atoms $a$ and $b$ respectively.
Now the reduced density matrix $\hat{\rho}_{a}$ for atom $a$ is
@@
$\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{4}-\Big{\langle}-\frac{1}{2}\Big{|}c_{3}\Big{]}$ (44)
@@
$\hat{J}_{a_{x^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\cos\theta_{1}-\hat{J}_{a_{z}}\sin\theta_{1}$ (48)
$\hat{J}_{a_{y^{\prime}}}$ $=$ $\hat{J}_{a_{y}}$ (49)
$\hat{J}_{a_{z^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\sin\theta_{1}+\hat{J}_{a_{z}}\cos\theta_{1},$ (50)
@@
It can be checked that
$\langle\hat{J}_{a_{x^{\prime}}}\rangle$ $=$ 00 0 (52)
$\langle\hat{J}_{a_{y^{\prime}}}\rangle$ $=$ $0,$ (53)
and hence the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ is now along the $z^{\prime}$-axis. Therefore,
@@
Similarly, it can be shown that for the atom $b$ the uncertainties in $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ are
@@
Now, the von Neumann entropy of the composite state in Eq. (12) is zero as it is a pure state. Therefore,
$S_{a,b}=0.$ (64)
But, the von Neumann entropies of the atoms $a$ and $b$, which are entangled are
@@
We now plot the uncertainties in Eqs. (42) and (43) and the von Neumann entropy in Eq. (65) with respect to the constant $c_{1}$ in FIG. 1.
Figure 1: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the atoms $a$ and $b$ with respect to the constant $c_{1}$. Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. $c_{1}$ has been plotted along horizontal axis and $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}$ and $S_{b}$ along the vertical axis. The thin line, thick line and the dotted line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$ and $S_{a}=S_{b}$ respectively.
We observe that though the von Neumann entropies $S_{a}$ and $S_{b}$ are always greater than 00, that is that of the composite state, in the whole range $0.001\leq c_{1}\leq 0.999$, as represented by the dotted line in FIG. 1, but both the uncertainties $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$ are greater than $\frac{1}{2}$ in the range $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$. The above ranges of $c_{1}$ correspond to the ranges of $c_{2}$ as $0.999\geq c_{2}\geq 0.966$ and $0.258\geq c_{2}\geq 0.001$. Therefore, both the uncertainties of the composite state in Eq. (12) are greater than those of the individual entangled atoms which have been represented in Eqs. (60) to (63). This is in contradiction with the prevailing consensus that more is the entropy, more is the uncertainty and more is the disorder.
@@
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{a}},$ (68)
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{b}},$ (69)
@@
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{a}},$ (70)
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{b}},$ (71)
@@
We also observe from FIG. 1 that when the von Neumann entropies $S_{a}=S_{b}$ of the individual atoms attain their peak values, the uncertainty $\Delta J_{x^{\prime}}$ of the composite state attains its peak value while the uncertainty $\Delta J_{y^{\prime}}$ attains $100\%$ squeezing.

### v2

@@ I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
We now calculate $\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle$, $\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle$, and $\langle\psi_{ab}|\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}|\psi_{ab}\rangle$ for the quantum state in Eq. (I.2), using the expression for $|\psi_{ab}\rangle$ given in Eq. (13).
Now,
$\hat{J}_{x}=\frac{1}{2}\Big{(}\hat{J}_{+}+\hat{J}_{-}\Big{)}.$
Therefore,
$\hat{J}_{x}^{2}=\frac{1}{4}\Big{(}\hat{J}_{+}^{2}+\hat{J}_{-}^{2}+\hat{{J}_{+}}\hat{{J}_{-}}+\hat{{J}_{-}}\hat{{J}_{+}}\Big{)},$ (22)
and
$\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}=\frac{1}{2}\Big{(}\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}+$
$\hat{J}_{-}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{-}\Big{)}.$ (23)
Since we are dealing here only with two atoms, $a$ and $b$, we have
$\hat{J}_{+}$ $=$ $\hat{J}_{a_{+}}+\hat{J}_{b_{+}},$
$\hat{J}_{-}$ $=$ $\hat{J}_{a_{-}}+\hat{J}_{b_{-}},$
$\hat{J}_{z}$ $=$ $\hat{J}_{a_{z}}+\hat{J}_{b_{z}}.$ (24)
Therefore,
$\hat{J}_{+}^{2}=\hat{J}_{a_{+}}^{2}+\hat{J}_{b_{+}}^{2}+2\hat{J}_{a_{+}}\hat{J}_{b_{+}},$ (25)
$\hat{J}_{+}\hat{J}_{-}+\hat{J}_{-}\hat{J}_{+}=\hat{J}_{a_{+}}\hat{J}_{a_{-}}+\hat{J}_{a_{-}}\hat{J}_{a_{+}}$
$+\hat{J}_{b_{+}}\hat{J}_{b_{-}}+\hat{J}_{b_{-}}\hat{J}_{b_{+}}+2\hat{J}_{a_{+}}\hat{J}_{b_{-}}+2\hat{J}_{a_{-}}\hat{J}_{b_{+}},$
and
$\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}=\hat{J}_{a_{+}}\hat{J}_{a_{z}}+\hat{J}_{a_{z}}\hat{J}_{a_{+}}+$
$\hat{J}_{b_{+}}\hat{J}_{b_{z}}+\hat{J}_{b_{z}}\hat{J}_{b_{+}}+2\Big{(}\hat{J}_{a_{+}}\hat{J}_{b_{z}}+\hat{J}_{a_{z}}\hat{J}_{b_{+}}\Big{)}.$
Now, as $A$, $B$, $C$, and $D$ are real, we have, using Eqs. (13) and (25),
$\langle\psi_{ab}|\hat{J}_{+}^{2}|\psi_{ab}\rangle=\bigg{[}\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}A+$
$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}B+\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}C+$
$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}D\bigg{]}\bigg{|}\Big{(}\hat{J}_{a_{+}}^{2}+\hat{J}_{b_{+}}^{2}+$
$2\hat{J}_{a_{+}}\hat{J}_{b_{+}}\Big{)}\bigg{|}\bigg{[}A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}$
$+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}\bigg{]}$
$=2AD.$ (28)
So,
$\langle\psi_{ab}|\hat{J}_{-}^{2}|\psi_{ab}\rangle=\langle\psi_{ab}|\hat{J}_{+}^{2}|\psi_{ab}\rangle^{*}$
$=2AD.$ (29)
Using Eqs. (13) and (26), we obtain
$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{-}+\hat{J}_{-}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle=$
$\bigg{[}\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}A+\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}B+$
$\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}C+\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}D\bigg{]}\bigg{|}$
$\Big{(}\hat{J}_{a_{+}}\hat{J}_{a_{-}}+\hat{J}_{a_{-}}\hat{J}_{a_{+}}+\hat{J}_{b_{+}}\hat{J}_{b_{-}}+\hat{J}_{b_{-}}\hat{J}_{b_{+}}+$
$2\hat{J}_{a_{+}}\hat{J}_{b_{-}}+2\hat{J}_{a_{-}}\hat{J}_{b_{+}}\Big{)}\bigg{|}\bigg{[}A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}$
$+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}\bigg{]}$
$=2\Big{(}A^{2}+B^{2}+C^{2}+D^{2}+2BC\Big{)}.$ (30)
Now, by virtue of Eq. (12), Eq. (30) reduces to
$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{-}+\hat{J}_{-}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle=2\Big{(}1+2BC\Big{)}.$
Therefore, using Eqs. (22), (28), (29), and (31), we get
$\langle\psi_{ab}|\hat{J}_{x}^{2}|\psi_{ab}\rangle=\frac{1}{2}\Big{(}1+2AD+2BC\Big{)}.$ (32)
Now, using Eqs. (13) and (24), we obtain
$\langle\psi_{ab}|\hat{J}_{z}^{2}|\psi_{ab}\rangle=\bigg{[}\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}A+$
$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}\frac{1}{2}\Big{|}B+\Big{\langle}\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}C+$
$\Big{\langle}-\frac{1}{2}\Big{|}\otimes\Big{\langle}-\frac{1}{2}\Big{|}D\bigg{]}\bigg{|}\Big{(}\hat{J}_{a_{z}}^{2}+\hat{J}_{b_{z}}^{2}+$
$2\hat{J}_{a_{z}}\hat{J}_{b_{z}}\Big{)}\bigg{|}\bigg{[}A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+$
$B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}$
$+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}\bigg{]}$
$=A^{2}+D^{2}.$ (33)
Similarly, using Eqs. (13) and (27), we get
$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle=\big{(}A-D\big{)}\times$
$\big{(}B+C\big{)}.$ (34)
Therefore,
$\langle\psi_{ab}|\Big{(}\hat{J}_{-}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{-}\Big{)}|\psi_{ab}\rangle=$
$\langle\psi_{ab}|\Big{(}\hat{J}_{+}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{+}\Big{)}|\psi_{ab}\rangle^{*}=$
$\big{(}A-D\big{)}\big{(}B+C\big{)}.$ (35)
So, using Eqs. (23), (34), and (35), we obtain
$\langle\psi_{ab}|\Big{(}\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\Big{)}|\psi_{ab}\rangle=$
$(A-D)(B+C).$ (36)
Therefore, using the expressions for $\langle\hat{J}_{x}\rangle$, $\langle\hat{J}_{z}\rangle$, $|\langle\hat{\mathbf{J}}\rangle|^{2}$, $\langle\hat{J}_{x}^{2}\rangle$, $\langle\hat{J}_{z}^{2}\rangle$, and $\langle\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\rangle$ from Eqs. (14), (16), (17), (32), (33), and (36), respectively, in Eq. (21), and simplifying, we get
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}^{2}|\psi_{ab}\rangle=\frac{1}{2}\big{(}A+D\big{)}^{2}.$ (37)
So, using Eqs. (3) and (37), we have
@@
Now,
$\hat{J}_{y}=\frac{1}{2i}\Big{(}\hat{J}_{+}-\hat{J}_{-}\Big{)}.$
Therefore,
$\langle\psi_{ab}|\hat{J}_{y}^{2}|\psi_{ab}\rangle=\frac{1}{4}\langle\psi_{ab}|\Big{(}\hat{{J}_{+}}\hat{{J}_{-}}+\hat{{J}_{-}}\hat{{J}_{+}}$
$-\hat{J}_{+}^{2}-\hat{J}_{-}^{2}\Big{)}|\psi_{ab}\rangle.$ (39)
So, using Eqs. (28), (29), and (31) in Eq. (39), and simplifying, we obtain
@@
These are the coordinate frame-independent inherent uncertainties of the composite state $|\psi_{ab}\rangle$ of the system.
We now proceed to calculate similar uncertainties in the atomic operators $\hat{J}_{a_{x^{\prime}}}$, $\hat{J}_{a_{y^{\prime}}}$, $\hat{J}_{b_{x^{\prime}}}$, and $\hat{J}_{b_{y^{\prime}}}$ of atoms $a$ and $b$, respectively.
Now, the reduced density matrix $\hat{\rho}_{a}$ for atom $a$ is
@@
$\Big{[}\Big{\langle}\frac{1}{2}\Big{|}c_{4}-\Big{\langle}-\frac{1}{2}\Big{|}c_{3}\Big{]}.$
@@
$\hat{J}_{a_{x^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\cos\theta_{1}-\hat{J}_{a_{z}}\sin\theta_{1},$ (47)
$\hat{J}_{a_{y^{\prime}}}$ $=$ $\hat{J}_{a_{y}},$ (48)
$\hat{J}_{a_{z^{\prime}}}$ $=$ $\hat{J}_{a_{x}}\sin\theta_{1}+\hat{J}_{a_{z}}\cos\theta_{1},$
@@
It can be verified, using Eqs. (45), (47), (48), and (49), that
$\langle\hat{J}_{a_{x^{\prime}}}\rangle$ $=$ $0,$
$\langle\hat{J}_{a_{y^{\prime}}}\rangle$ $=$ $0,$
and hence, the mean spin vector $\langle\hat{\mathbf{J}}_{a}\rangle$ is now along the $z^{\prime}$-axis. Therefore,
@@
Similarly, it can be shown that, for atom $b$, the uncertainties in $\hat{J}_{b_{x^{\prime}}}$ and $\hat{J}_{b_{y^{\prime}}}$ are
@@
Now, the von Neumann entropy of the composite state in Eq. (I.2) is zero, as it is a pure state. Therefore,
$S_{a,b}=0.$
But the von Neumann entropies of atoms $a$ and $b$, which are entangled, are
@@
We now plot the uncertainties in Eqs. (42) and (43), and the von Neumann entropy in Eq. (60), with respect to the constant $c_{1}$ in FIG. 1.
Figure 1: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ of atoms $a$ and $b$ with respect to the constant $c_{1}$. Here, we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. The constant $c_{1}$ is plotted along the horizontal axis, while $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the entropies $S_{a}$ and $S_{b}$ are plotted along the vertical axis. The thin line, thick line, and dotted line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and $S_{a}=S_{b}$, respectively.
We observe that, although the von Neumann entropies $S_{a}$ and $S_{b}$ are always greater than 00 —that is, greater than that of the composite state —over the entire range $0.001\leq c_{1}\leq 0.999$, as represented by the dotted line in FIG. 1, both the uncertainties $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$ are greater than $\frac{1}{2}$ in the ranges $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$. The above ranges of $c_{1}$ correspond to the ranges of $c_{2}$ as $0.999\geq c_{2}\geq 0.966$, and $0.259\geq c_{2}\geq 0.045$, respectively. Therefore, both the uncertainties of the composite state in Eq. (I.2) are greater than those of the individual entangled atoms, which have been represented in Eqs. (56) to (59). This is in contradiction with the prevailing idea that the more the entropy, the greater the uncertainty and disorder.
@@
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{a}},$
$\Delta J_{x^{\prime}_{a,b}}>\Delta J_{x^{\prime}_{b}},$
@@
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{a}},$
$\Delta J_{y^{\prime}_{a,b}}>\Delta J_{y^{\prime}_{b}},$
@@
We also observe from FIG. 1 that, when the von Neumann entropies $S_{a}=S_{b}$ of the individual atoms attain their peak values, the uncertainty $\Delta J_{x^{\prime}}$ of the composite state attains its peak value, while the uncertainty $\Delta J_{y^{\prime}}$ attains $100\%$ squeezing.
Figure 2: Plot of the uncertainties $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ of atoms $a$ and $b$ with respect to the constant $c_{1}$. Here, we have taken $c_{3}=\frac{3}{8}$, $c_{4}=\frac{\sqrt{55}}{8}$, $c_{5}=\frac{4}{7}$ and $c_{6}=\frac{\sqrt{33}}{7}$. The constant $c_{1}$ is plotted along the horizontal axis, while $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the entropies $S_{a}$ and $S_{b}$ are plotted along the vertical axis. The thick line, dashed line, and thin line represent $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and $S_{a}=S_{b}$, respectively.
In FIG. 1, we have chosen the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$ as $\frac{1}{\sqrt{2}}$; that is, the quantum state $|\psi_{ab}\rangle$ is symmetric with respect to these parameters. Now, it can be seen that the above-mentioned feature also holds for asymmetric combinations of these parameters. This is because we observe from Eqs. (42) and (43) that while $\Delta J_{x^{\prime}}$ depends on the parameters $c_{1},c_{2},c_{3},c_{4},c_{5}$, and $c_{6}$, $\Delta J_{y^{\prime}}$ depends only on $c_{1}$ and $c_{2}$. Therefore, variations in the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$ do not affect $\Delta J_{y^{\prime}}$, but they do affect $\Delta J_{x^{\prime}}$. Thus, $\Delta J_{y^{\prime}}$ is always greater than that of the individual entangled atoms in the ranges $0.001\leq c_{1}\leq 0.258$ and $0.966\leq c_{1}\leq 0.999$. However, whether $\Delta J_{x^{\prime}}$ will be greater than that of the individual entangled atoms in the above-mentioned ranges of $c_{1}$ depends on the choice of the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$. Now, in Eq. (42), the minimum value of the factor $\sqrt{\Big{(}\frac{1}{2}+c_{1}c_{2}\Big{)}}$ is $0.707$. Therefore, $\Delta J_{x^{\prime}}$ is greater than $\frac{1}{2}$ —that is, greater than that of the individual enatngled atoms —as long as the factor $(c_{3}c_{5}+c_{4}c_{6})$ in Eq. (42) is greater than $0.707$. Thus, it does not matter whether the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$ are symmetric or asymmetric, as long as their combination makes the factor $(c_{3}c_{5}+c_{4}c_{6})$ in Eq. (42) greater than $0.707$. As an example, in FIG. 2, we plot $\Delta J_{x^{\prime}}$, $\Delta J_{y^{\prime}}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ for an asymmetric combination of the parameters $c_{3},c_{4},c_{5}$, and $c_{6}$, where $c_{3}=\frac{3}{8}$, $c_{4}=\frac{\sqrt{55}}{8}$, $c_{5}=\frac{4}{7}$, and $c_{6}=\frac{\sqrt{33}}{7}$. We observe the same result: that both $\Delta J_{x^{\prime}}$ and $\Delta J_{y^{\prime}}$, for the composite state, are greater than those of the individual entangled atoms in the same ranges of the parameters $c_{1}$ and $c_{2}$, as mentioned previously.